# TerraMind: Large-Scale Generative Multimodality for Earth Observation

### v1

@@ Abstract
We present TerraMind, the first any-to-any generative, multimodal foundation model for Earth observation (EO). Unlike other multimodal models, TerraMind is pretrained on dual-scale representations combining both token-level and pixel-level data across modalities. On a token level, TerraMind encodes high-level contextual information to learn cross-modal relationships, while on a pixel level, TerraMind leverages fine-grained representations to capture critical spatial nuances. We pretrained TerraMind on nine geospatial modalities of a global, large-scale dataset. In this paper, we demonstrate that (i) TerraMind’s dual-scale early fusion approach unlocks a range of zero-shot and few-shot applications for Earth observation, (ii) TerraMind introduces “thinking in modalities” (TiM)—the capability of generating additional artificial data during finetuning and inference to improve the model output—and (iii) TerraMind achieves beyond state-of-the-art performance in community-standard benchmarks for EO like PANGAEA. The pretraining dataset, the model weights, and our code will be open-sourced under a permissive license.
@@ 1 Introduction
Figure 1: TerraMind represents the first any-to-any generative, and large-scale multimodal model for Earth observation pre-trained on 1 trillion tokens from global geospatial data. The model digests dual-scale representations at pixel-level and token-level simultaneously. TerraMindv1 outperforms other deep learning models for Earth observation in downstream applications and unlocks (i) generation, (ii) zero-shot and finetuning applications, and (iii) ”Thinking-in-Modalities” finetuning and inference.
In this work, we introduce TerraMind, the first large-scale, any-to-any generative, multimodal model for EO. With TerraMind, we introduce a dual-scale pretraining on pixel-level and token-level for generative multi-modality extending previous approaches like 4M [52] and demonstrate benefits over training primarily on tokens. TerraMind encodes high-level contextual information in tokens to enable correlation learning and scaling, while, additionally capturing important fine-grained representations using pixel-level inputs. During the pretraining, TerraMind predicts masked target tokens so that our pretraining objective boils down to a cross-modal patch classification problem that results in high-quality latent spaces. TerraMind is pretrained on a custom global-scale geospatial dataset named TerraMesh with 9 million samples that have been aligned spatiotemporally and across modalities [7]. In addition to radar and optical satellite imagery from the Copernicus Sentinel-1 (S-1) and Sentinel-2 (S-2) missions, our dataset contains task-specific modalities such as land use/land cover (LULC) and normalized difference vegetation index (NDVI) maps, metadata like digital elevation models (DEM) and geographic coordinates, and natural language in the form of captions. To the best of our knowledge, TerraMind represents the first truly generative, multimodal deep learning model for EO. We provide an overview on TerraMind’s performance in the community-standard benchmark PANGAEA [49] in Figure 2 and highlight the generative capabilities of TerraMind in Figure 3. In summary, our key contributions are as follows: (i) We introduce a dual-scale approach for generative multimodal pre-training leveraging data on pixel-level and token-level, which outperforms other fusion approaches, enables generation of artificial data, and facilitates zero-shot and few-shot applications. (ii) We introduce thinking in modalities – similar to chain-of-thought approaches in LLMs – for multi-modal models in EO, demonstrating that infusing generated data during finetuning improves the performance. (iii) We demonstrate that TerraMind outperforms other geospatial foundation models both in unimodal and multimodal settings.
@@ 2 Related Work
Multimodality in EO. Multimodality in EO originates from data fusion and is typically understood as the integration of SAR and optical data [13, 28, 38, 25] or the combination of optical data with vector data [5]. Some studies have explored alternative combinations of data. In [15], the authors introduce a contrastive framework for comparing RS images and street views. Even different optical sensors can be considered different modalities [61, 48]. Similarly, several multi-view images (i.e. multimodal) datasets [44, 26, 54] are introduced. More recent approaches combined text and images [40], both for discriminative [42] and generative [34] purposes. Lately, different GFMs are trained in a multimodal way [4, 54, 68], still focusing either on a specific set of modalities (e.g., vision [68]) or tasks (e.g., generative [34]). To the best of our knowledge, no existing model has combined a wide and diverse amount of modalities both for discriminative and generative purposes, as TerraMind does. In Table 1, we provide a comparison of TerraMind with related model architectures.
@@ 4.1 Tokenization
Image-like modalities. We train autoencoder-based architectures with a quantization step in the bottleneck for image-like modalities such as Sentinel-1, Sentinel-2, LULC, NDVI, and DEM. Tokenizer encoders process an input image and generate a latent representation for each 16$\times$16 patch, which is then discretized with finite-scalar-quantization (FSQ) [51] into one of $N$ codewords, where $N$ represents the codebook size. All tokenizers use a vocabulary size of 16K besides the simpler LULC modality for which we use 4K. These codewords are then used by the diffusion decoder to reconstruct the original image. By mapping each codeword to a unique integer in $\{0,1,\dots,N-1\}$, we obtain discrete tokens for each image patch. We pretrain the tokenizers in a self-supervised setting. FSQ as quantization method enhances training stability [51] compared to vector quantization [63]. FSQ improves training stability by using a fixed codebook instead of a learnable one, eliminating the need for codebook-related loss terms. Notably, FSQ is heavily influenced by ideas of neural compression [27]. For our 12 channel Sentinel2-L2A data, we achieve compression rates of over 3000x by quantizing the data. We summarize the architecture of our tokenizers in Figure 4.
@@ 4.2 Pre-training
Architecture. TerraMind uses a symmetric Transformer-based encoder-decoder architecture proposed in [52], which is designed to process sequences of multimodal tokens. In addition to discrete tokens, our models also accept cross-modal pixel-level inputs, specifically satellite imagery and digital elevation maps. In contrast to [52], we leverage pixel-level data across a range of input modalities to introduce dual-scale feature representations in the input space, which help to structure the embedding space. For pixel-level inputs, we apply learnable patch-wise linear projections to generate patch embeddings for each 16$\times$16 patch, similar to the approach used in ViT [17]. We illustrate the pretraining paradigm in Figure 5 with an example of four modalities. The model is agnostic to processing tokens or patches in the input space, while the target is generally token-level data. In total, we use six pixel-level modalities and eight token-level modalities. Some of those modalities overlap, i.e. they are present both as pixel-level and token-level information which allows the model to learn a cross-modal patch-wise classification in one of the $N$ codebook classes.
@@ 4.3 Generation
After pretraining, TerraMind is able to generate tokens for any modality, conditioned on any subset of input modalities. The generative capabilities also unlock various zero-shot tasks, such as water body segmentation, vegetation assessments, land use segmentation, image captioning, geolocation prediction, and text-to-image generation. For the generation of image-like modalities, the decoder receives mask tokens for the modality to be generated and predicts the corresponding tokens based on the processed input from the encoder. For sequence-like modalities, the decoder is initialized with a start-of-sequence token and generates the output autoregressively. After generating all output tokens, the corresponding tokenizer allows to map from token-space to image or text space. TerraMind additionally supports chained generation which ensures consistency across generated modalities. The chained generation represents a conditional probability distribution where the prior probability distribution $\mathbf{P(M_{1})}$ is given by the input modality $\mathbf{M_{1}}$, and all subsequent modalities $\mathbf{M_{2}}$ to $\mathbf{M_{K}}$ are generated based on their conditioning on the input modality and potential additional ones:
@@ 5.1 Foundational experiments
Multimodality vs. unimodality. As a first motivational experiment, we outline the benefit of leveraging multimodal data in Earth observation at the example of water body mapping. For this, we leverage the ViT-B encoders from the unimodal tokenizer models for S-1, S-2, and LULC. For fine-tuning, we concatenate the embeddings of the encoders, and train a segmentation head with four ConvNeXt [43] blocks (late fusion). The results in Table 2 (left) suggest that regardless of which modalities we combine, the combination of two modalities always outperforms each unimodal model. Combining all three modalities achieves the best overall performance.
@@ 5.2 Generation experiments
TerraMind supports the generation of any modality from any other modality. In the following, we provide examples of the generation performance starting from: (i) an information-rich modality, like optical S-2 L2A data, and (ii) minimal information based on the geolocation. In Figure 3, we observe that TerraMind performs very well in generating image-like modalities like radar data, land use maps, and digital elevation maps from optical S-2 L2A data. Interestingly, TerraMind is also able to geolocate the optical image with a low deviation of less than 0.5 degree latitude/longitude. We provide additional experiments on the accuracy of geo-localization prediction in Section 5.3.
@@ 5.3 Zero-shot experiments
Geo-localization. TerraMind has strong capabilities in predicting the geolocation of a specific data instance. To better visualize the geo-localization capabilities, we prompt the model for the most likely locations of the land use class “bare land” (deserts etc.) in a Monte-Carlo-sampling in Figure 7(a). The probability distribution of the model fits the expectation of where to find bare land accurately, highlighting the Sahara region and middle-east, as well as Mexico and Southern California.
(a) Prediction distribution of the land use class “bare land” with a sampling temperature of $T=1.0$. TerraMind has an accurate internal representation of the geolocation of specific contexts, like land use classes.
(b) Distribution of predicted geo-locations for an optical S-2 L2A sample from the validation set. $\bullet$ is the correct location, $\bullet$ are Monte-Carlo sampled locations from TerraMind, $\bullet$ represents the distribution of training locations. TerraMind’s geo-localization seems to be based on similar optical samples in the training dataset for which TerraMind then outputs the geolocation.
Figure 7: Monte-Carlo sampling approach for TerraMind’s geo-location prediction.
To better understand how TerraMind assigns geolocations to samples, we further employ a Monte-Carlo sampling on the latitude-longitude grid for an optical tile from the validation data in Figure 7(b). We observe that while TerraMind is not predicting the correct geolocation ($\bullet$), there is a very high likelihood that the predicted geolocation is one of the adjacent grid points ($\bullet$). Thus, TerraMind likely looks out for similar looking optical samples in the training data ($\bullet$) and returns their geolocation.
@@ 5.4 Few-shot experiments
CLIP-ViT-B/16 S-2 RGB 52.39 28.13
CLIP-ViT-B/16 NAIP 31.73
@@ 5.5 Fine-tuning experiments
Besides the novel capabilities that TerraMind introduces, we benchmark the fine-tuning performance of TerraMind in both unimodal and multimodal settings following the community-standard PANGAEA benchmark [49]. We summarize the results in Table 5. Overall, TerraMindv1-B outperforms all other GeoFMs by at least 3pp avg. mIoU. Importantly, we report additional results in Table 11 of the supplementary material that show how TerraMind is the only foundation model approach in EO that outperforms task-specific U-Net models across the PANGAEA benchmark. Performance is approximately 2pp avg. mIoU higher for Terramindv1-L, with a peak of 5pp in multimodal datasets. Furthermore, TerraMindv1-L outperforms also specialised ViT baselines by 5pp avg. mIoU. Finally, TerraMindv1-B outperforms TerraMindv1-B-single by 1pp. avg. mIoU. The latter result demonstrates the benefit of using dual-scale features when comparing v1-B-single with v1-B. Note that per suggestion of the PANGAEA authors, we exclude the xView2 and BioMassters task as we could not reproduce the reported performances.
Model BurnSr MADOS PASTIS Sen1Fl11 FBP DEN CTM-SS SN7 AI4Farms Avg. mIoU $\uparrow$ Avg. Rank $\downarrow$ CROMA 82.42 67.55 32.32 90.89 51.83 38.29 49.38 59.28 25.65 55.29 5.06 DOFA 80.63 59.58 30.02 89.37 43.18 39.29 51.33 61.84 27.07 53.59 6.33 GFM-Swin 76.90 64.71 21.24 72.60 67.18 34.09 46.98 60.89 27.19 52.42 8.00 Prithvi 1.0 100M 83.62 49.98 33.93 90.37 46.81 27.86 43.07 56.54 26.86 51.00 8.89 RemoteCLIP 76.59 60.00 18.23 74.26 69.19 31.78 52.05 57.76 25.12 51.66 9.22 SatlasNet 79.96 55.86 17.51 90.30 50.97 36.31 46.97 61.88 25.13 51.65 8.56 Scale-MAE 76.68 57.32 24.55 74.13 67.19 35.11 25.42 62.96 21.47 49.43 9.44 SpectralGPT 80.47 57.99 35.44 89.07 33.42 37.85 46.95 58.86 26.75 51.87 8.11 S.-S12-MoCo 81.58 51.76 34.49 89.26 53.02 35.44 48.58 57.64 25.38 53.02 8.00 S.-S12-DINO 81.72 49.37 36.18 88.61 51.15 34.81 48.66 56.47 25.62 52.51 8.67 S.-S12-MAE 81.91 49.90 32.03 87.79 51.92 34.08 45.80 57.13 24.69 51.69 10.28 S.-S12-Data2Vec 81.91 44.36 34.32 88.15 48.82 35.90 54.03 58.23 24.23 52.22 8.72 TerraMindv1-B 82.42 69.52 40.51 90.62 59.72 37.87 55.80 60.61 28.12 58.35 2.94 TerraMindv1-L 82.93 75.57 43.13 90.78 63.38 37.89 55.04 59.98 27.47 59.57 2.56
Table 5: Performance evaluation of frozen TerraMind encoder with trainable UperNet head following the PANGAEA evaluation protocol across nine benchmark datasets. We report the mIoU (%) $\uparrow$ and highlight the best model in bold and the second best is underscored.
Finally, we assess the impact of leveraging multimodal data as input to TerraMindv1-B compared to utilizing either optical or radar data as unimodal input to better understand the effect of leveraging multimodal data in finetuning. We observe that across all three multimodal tasks, TerraMindv1-B performs best with access to both optical and radar data.

### v2

@@ Abstract
We present TerraMind, the first any-to-any generative, multimodal foundation model for Earth observation (EO). Unlike other multimodal models, TerraMind is pretrained on dual-scale representations combining both token-level and pixel-level data across modalities. On a token level, TerraMind encodes high-level contextual information to learn cross-modal relationships, while on a pixel level, TerraMind leverages fine-grained representations to capture critical spatial nuances. We pretrained TerraMind on nine geospatial modalities of a global, large-scale dataset. In this paper, we demonstrate that (i) TerraMind’s dual-scale early fusion approach unlocks a range of zero-shot and few-shot applications for Earth observation, (ii) TerraMind introduces “thinking in modalities” (TiM)—the capability of generating additional artificial data during finetuning and inference to improve the model output—and (iii) TerraMind achieves beyond state-of-the-art performance in community-standard benchmarks for EO like PANGAEA. The pretraining dataset, the model weights, and our code have been open-sourced under a permissive license.
Figure 1: TerraMind represents the first any-to-any generative, and large-scale multimodal model for Earth observation pre-trained on 1 trillion tokens from global geospatial data. The model digests multi-scale representations at pixel-level and token-level simultaneously. TerraMindv1 outperforms other deep learning models for Earth observation in downstream applications and unlocks (i) generation, (ii) zero-shot and finetuning applications, and (iii) ”Thinking-in-Modalities” finetuning and inference.
@@ 1 Introduction
In this work, we introduce TerraMind, the first large-scale, any-to-any generative, multimodal model for EO. With TerraMind, we introduce a dual-scale pretraining on pixel-level and token-level for generative multi-modality extending previous approaches like 4M [52] and demonstrate benefits over training primarily on tokens. TerraMind encodes high-level contextual information in tokens to enable correlation learning and scaling, while, additionally capturing important fine-grained representations using pixel-level inputs. During the pretraining, TerraMind predicts masked target tokens so that our pretraining objective boils down to a cross-modal patch classification problem that results in high-quality latent spaces. TerraMind is pretrained on a custom global-scale geospatial dataset named TerraMesh with 9 million samples that have been aligned spatiotemporally and across modalities [7]. In addition to radar and optical satellite imagery from the Copernicus Sentinel-1 (S-1) and Sentinel-2 (S-2) missions, our dataset contains task-specific modalities such as land use/land cover (LULC) and normalized difference vegetation index (NDVI) maps, metadata like digital elevation models (DEM) and geographic coordinates, and natural language in the form of captions. To the best of our knowledge, TerraMind represents the first truly generative, multimodal deep learning model for EO. Furthermore, in contrast to many current remote sensing models that utilize masked autoencoders (MAE) like [54], contrastive learning (CL), or diffusion techniques, TerraMind uniquely demonstrates benefits of leveraging the autoregressive paradigm for Earth observation.
We provide an overview on TerraMind’s performance in the community-standard benchmark PANGAEA [49] in Figure 2 and highlight the generative capabilities of TerraMind in Figure 3. In summary, our key contributions are as follows: (i) We introduce a dual-scale approach for generative multimodal pre-training leveraging data on pixel-level and token-level, which outperforms other fusion approaches, enables generation of artificial data, and facilitates zero-shot and few-shot applications. (ii) We introduce thinking in modalities – similar to chain-of-thought approaches in LLMs – for multi-modal models in EO, demonstrating that infusing generated data during finetuning improves the performance. (iii) We demonstrate that TerraMind outperforms other geospatial foundation models both in unimodal and multimodal settings.
@@ 2 Related Work
Multimodality in EO. Multimodality in EO originates from data fusion and is typically understood as the integration of SAR and optical data [13, 28, 38, 25] or the combination of optical data with vector data [5]. Some studies have explored alternative combinations of data. In [15], the authors introduce a contrastive framework for comparing RS images and street views. Even different optical sensors can be considered different modalities [61, 48]. Similarly, several multi-view images (i.e. multimodal) datasets [44, 26, 54] are introduced. More recent approaches combined text and images [40], both for discriminative [42] and generative [34] purposes. Lately, different GFMs are trained in a multimodal way [4, 54, 68], still focusing either on a specific set of modalities (e.g., vision [68], [3]) or tasks (e.g., generative [34]). Compared to multi-scale high-quality generation models for optical data, like MetaEarth [71], our approach allows to generate any modality from any other pretraining modality. To the best of our knowledge, no existing model has combined a wide and diverse amount of modalities both for discriminative and generative purposes, as TerraMind does. In Table 1, we provide a comparison of TerraMind with related model architectures.
@@ 4.1 Tokenization
Image-like modalities. We train autoencoder-based architectures with a quantization step in the bottleneck for image-like modalities such as Sentinel-1, Sentinel-2, LULC, NDVI, and DEM. Tokenizer encoders process an input image and generate a latent representation for each 16$\times$16 patch, which is then discretized with finite-scalar-quantization (FSQ) [51] into one of $N$ codewords, where $N$ represents the codebook size. All tokenizers use a vocabulary size of 16K besides the simpler LULC modality for which we use 4K. These codewords are then used by the diffusion decoder to reconstruct the original image. The benefit of leveraging diffusion decoders lies in facilitating cross-modal generation in TerraMind by transforming tokens back into images. By mapping each codeword to a unique integer in $\{0,1,\dots,N-1\}$, we obtain discrete tokens for each image patch. We pretrain the tokenizers in a self-supervised setting. FSQ as quantization method enhances training stability [51] compared to vector quantization [63]. FSQ improves training stability by using a fixed codebook instead of a learnable one, eliminating the need for codebook-related loss terms. Notably, FSQ is heavily influenced by ideas of neural compression [27]. For our 12 channel Sentinel2-L2A data, we achieve compression rates of over 3000x by quantizing the data. We summarize the architecture of our tokenizers in Figure 4. The main objective of the overall tokenizer is to encode image patches consistently into discrete tokens based on semantic similarity to enable cross-modal correlation learning. Thus, the loss of certain details is an expected trade-off, since the focus is on grouping similar patches rather than preserving all fine-grained features. Naturally, more accurate reconstructions faciliate the cross-modal generation, which is an important component of TerraMind – the main focus however lies on consistent cross-modal correlation learning. We provided further details on the pretraining of the tokenizers in the supplementary material and open-source code and weights for the tokenizer models.
@@ 4.2 Pre-training
Architecture. TerraMind uses a symmetric Transformer-based encoder-decoder architecture proposed in [52], which is designed to process sequences of multimodal tokens. In addition to discrete tokens, TerraMind accepts pixel-level inputs, specifically satellite imagery and digital elevation maps. For pixel-level inputs, we apply learnable patch-wise linear projections to generate patch embeddings for each 16$\times$16 patch, similar to the approach used in ViT [17].
Dual-scale early fusion. In contrast to [52], we not only embed token-level data but additionally leverage pixel-level data across a range of input modalities to introduce a dual-scale feature representation to support the structuring of the embedding space. Both tokens and patches represent a 16x16 pixel area. Tokens represent this area via a single discrete integer value, while the image patches describe the same area with the actual floating point sensor data. During pretraining, the model, thus, not only learns a correlation between modalities (i.e., cross-modal learning) but also between different levels of abstraction within the same modality. The low-level token information enables the cross-modal correlation learning, while adding pixel level input accounts for spatial nuances. Based on dual-scale features the model further learns to better structure pixel-level data in the embedding space via the corresponding information from the discrete token. We illustrate the pretraining paradigm in Figure 5. The model is agnostic to processing tokens or patches in the input space, while the target is generally token-level data. We use six pixel-level modalities and eight token-level modalities. Optical, radar, and elevation data is present both in pixel-space and token-space.
@@ 4.3 Generation
After pretraining, TerraMind can generate tokens for any modality, conditioned on any subset of input modalities. The generative capabilities further unlock various zero-shot tasks, such as water body segmentation, land use segmentation, image captioning, and geolocation prediction. For the generation of image-like modalities, the decoder receives mask tokens for the modality to be generated and predicts the corresponding tokens based on the encoded input. For sequence-like modalities, the decoder generates the output autoregressively. After generating tokens from the target modality, the corresponding tokenizer decoder allows to map from token-space to image or text space.TerraMind further supports chained generation which ensures consistency across generated modalities. The chained generation represents a conditional probability distribution where the prior probability distribution $\mathbf{P(M_{1})}$ is given by the input modality $\mathbf{M_{1}}$, and all subsequent modalities $\mathbf{M_{2}}$ to $\mathbf{M_{K}}$ are generated conditioned on the input modality and potentially other generated modalities:
@@
4.4 Thinking-in-Modalities
Thinking in Modalities (TiM) is a recursive fine-tuning and inference technique designed to enhance multimodal learning by leveraging the generative capabilities of the model itself. Given an input $x\in\mathcal{X}$ (e.g., an optical satellite image), the model first generates additional synthetic modalities $\tilde{x}=f_{\text{gen}}(x)$ on a token-level using a learned generative function $f_{\text{gen}}$. These generated tokens are then concatenated with the original input and jointly processed by the downstream model $f$ (e.g., TerraMind encoder with a segmentation head), yielding the final output $y=f(x,f_{\text{gen}}(x))$. This formulation allows the model to reason over both observed and inferred modalities, effectively enriching the input space. TiM can leverage multiple generated modalities which are then generated in a chained approach (see Eq. 4). For example, for $k$ modalities, the input is augmented with newly generated modalities:
$\tilde{x}^{(k+1)}=\tilde{x}^{(k)}\cup f_{\text{gen}}(x^{(k)}),$ (5)
and the final model output is described by:
$y=f(\tilde{x}^{(K)}).$ (6)
This recursive augmentation mimics a chain-of-thought process, enabling the model to iteratively refine its internal representation and improve performance on downstream tasks, particularly in scenarios with missing modalities.
@@ 5.1 Foundational experiments
Multimodality vs. unimodality. As a first motivational experiment, we outline the benefit of using multimodal data in Earth observation at the example of water body mapping. Specifically, we leverage the ViT-B encoders from the unimodal tokenizer models for S-1, S-2, and LULC, concatenate their embeddings, and train a segmentation head with four ConvNeXt [43] blocks as a late fusion approach. The results in Table 2 (left) suggest that regardless of which modalities we combine, the combination of two modalities always outperforms each unimodal model. Combining all three modalities achieves the best overall performance.
@@ 5.2 Generation experiments
(a) Input: S-2 L2A data capturing Singapore in January 2025.
(b) Generation: S-1 RTC composition generated by TerraMind.
(c) Input: S-2 L2A data capturing Nothern Spain in January 2025.
(d) Generation: S-1 GRD composition generated by TerraMind.
Figure 6: Generated Sentinel-1 data using TerraMind over Singapore and Northern Spain.
[…]
S-1 GRD $\rightarrow$ S-2 L2A 0.074 0.116 0.750 26.210
S-1 GRD $\rightarrow$ DEM 163.0 320.8 0.878 20.694
S-1 GRD $\rightarrow$ NDVI 0.180 0.225 0.438 18.990
S-1 RTC $\rightarrow$ S-2 L2A 0.113 0.194 0.695 24.251
S-1 RTC $\rightarrow$ DEM 298.8 799.2 0.873 20.009
S-1 RTC $\rightarrow$ NDVI 0.172 0.211 0.465 19.529
S-2 L2A $\rightarrow$ S-1 GRD 2.942 3.877 0.531 28.678
S-2 L2A $\rightarrow$ S-1 RTC 2.636 3.391 0.430 28.993
S-2 L2A $\rightarrow$ DEM 215.8 745.5 0.942 20.616
Table 3: Quantitative evaluation of generations on unseen global validation dataset using 10 diffusion steps. MAE and RMSE metrics are in physical units: meter (DEM), reflectance (S-2), and db (S-1).
TerraMind supports the generation of any modality from any other modality. In the following, we provide examples of the generation performance starting from: (i) an information-rich modality, like optical S-2 L2A data, and (ii) minimal information based on the geolocation. In Figure 3, we observe that TerraMind performs very well in generating image-like modalities like radar data, land use maps, and digital elevation maps from optical S-2 L2A data. We provide a quantitative overview on the quality of the generations on unseen validation data in Table 3. Overall, we observe an interesting asymmetry in the generative performance of TerraMind where (a) radar-to-optical generation achieves reasonable generation quality in terms of SSIM and PSNR – indicating structural and visual fidelity with some perceptual degradation – and (b) optical-to-radar generation yields higher PSNR values but lower SSIM, suggesting visually plausible outputs that lack strong structural alignment. The quality of generated DEM suggest to be structurally very strong, but noisy. The errors for DEM generations suggest that the generations structurally make sense but the level of altitude is difficult to infer for the model. We compare these scores with the reconstruction quality of the auto-encoding tokenizers in the supplementary material that can serve as upper bounds. Additionally, we provide experiments on the generation quality using token-level instead of pixel-level inputs in the supplementary material. Finally, we demonstrate the quality of large-scale generations across Singapore and Northern Spain in Figures 19 and 21.
@@ 5.3 Zero-shot experiments
Geo-localization. TerraMind is able to predict the geolocation of a specific data instance. To better visualize the geo-localization capabilities, we prompt the model for the most likely locations of the land use class “bare land” (deserts etc.) in a Monte-Carlo-sampling in Figure 7. The probability distribution of the model fits the expectation of where to find bare land, highlighting the Sahara region and middle-east, as well as Mexico and Southern California.
Figure 7: Prediction distribution of the land use class “bare land” with a sampling temperature of $T=1.0$. TerraMind has an accurate internal representation of the geolocation of specific contexts, like land use classes.
@@ 5.4 Few-shot experiments
CLIP-ViT-B/16 S-2 RGB 57.00 29.15
CLIP-ViT-B/16 NAIP 32.01
@@ 5.5 Fine-tuning experiments
Table 6: Performance evaluation of TerraMind across nine benchmark datasets using the PANGAEA evaluation protocol. Higher mIoU values ($\uparrow$) and lower rank values ($\downarrow$) are reported. The best model per column is highlighted in bold and the second best is underscored. We indicate unimodal datasets with , demonstrating that TerraMind outperforms other benchmarks in both unimodal and multimodal settings. All foundation model encoders are frozen, while U-Net and ViT baselines are trained from scratch for each specific task.
Besides the novel capabilities that TerraMind introduces, we benchmark the fine-tuning performance of TerraMind in both unimodal and multimodal settings following the community-standard PANGAEA benchmark [49]. We summarize the results in Table 6. Overall, TerraMindv1-B outperforms all other GeoFMs by at least 3pp avg. mIoU. Importantly, we observe that TerraMind is the only foundation model approach in EO that across the PANGAEA benchmark outperforms task-specific U-Net models. Performance is approximately 2pp avg. mIoU higher for TerraMindv1-L, with a peak of 5pp in multimodal datasets. Furthermore, TerraMindv1-L outperforms also specialised ViT baselines by 5pp avg. mIoU. Finally, TerraMindv1-B outperforms TerraMindv1-B-single by 1pp. avg. mIoU. The latter result demonstrates the benefit of using dual-scale features when comparing v1-B-single with v1-B. Note that per suggestion of the PANGAEA authors, we exclude the xView2 and BioMassters task as we could not reproduce the reported performances. Finally, we assess the impact of leveraging multimodal data as input to TerraMindv1-B compared to utilizing either optical or radar data as unimodal input to better understand the effect of leveraging multimodal data in finetuning. We observe that across all three multimodal tasks, TerraMindv1-B performs best with access to both optical and radar data with a significant improvement for CTM-SS.